# ELFATT: Efficient Linear Fast Attention for Vision Transformers

### v1

@@ III-D Efficient Linear Fast Attention Mechanism
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}\\ &\leq\left(\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|\textit{{U}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}+\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\right.\\ &\left.\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\right)\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$ (10)
@@
If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the total approximation error is bounded as follows,
@@
$\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}\\ &\leq\left(\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|\textit{{U}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}+\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\right.\\ &\left.\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\right)\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$ (11)
Eq. (8) denotes the single-head vanilla softmax-based attention mechanism approximated by using ELFATT, and Inequality (9) denotes the corresponding approximation error. From the right-hand side of Eq. (8), each ELFATT attention process consists of two heads. Hence, ELFATT can directly become an approximation of the double-head vanilla softmax-based attention mechanism as follows,
$\begin{split}&\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\right.\\ &\left.g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})\right)\right].\end{split}$ (12)
It is obvious to see that
$\begin{split}&\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},\right.\\ &\left.{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]=\left[\left({\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right)\bar{\textit{{V}}},\right.\\ &\left.\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\tilde{\textit{{V}}}\right].\end{split}$
@@
$\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},\right.\right.\\ &\left.\left.{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$ (13)
If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, we have
@@
$\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},\right.\right.\\ &\left.\left.{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\leq\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|\textit{{Z}}-\emph{{{U}}}_{m}\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$ (14)
If $\left|\frac{c}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, one has
@@
$\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},\right.\right.\\ &\left.\left.{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\leq\left|\frac{c}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|\textit{{Z}}-\emph{{{U}}}_{m}\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$ (15)
Inequalities (13-15) give a tighter bound than Inequalities (9-11).
Since each ELFATT attention block corresponds to two parallel heads, $s$ ELFATT will be needed for the approximation of $2\times s$ heads of vanilla softmax-based attention according to Eq. (12). For the approximation of $12\times s+$ 1 heads of vanilla softmax-based attention, $12\times s+$ 1 ELFATT will be needed according to Eq. (8).
III-E Positional Encoding
Different positional encoding mechanisms such as absolute positional encoding [38], relative positional encoding [39], and conditional positional encoding [40], have been proposed to make use of the ordering information of sequences in vision transformers [38, 23, 40]. Among these positional encoding mechanisms, the locally enhanced positional encoding (LePE) [23] mechanism shows more powerful local positional information enhancement and brings a higher performance gain for vision transformers. Hence, LePE is selected as the positional encoding mechanism for ELFATT. After introducing LePE, Eq. (8) will become as follows,
@@ III-E Positional Encoding
$\begin{split}&{\rm exp}\left(\textit{{Q}}\textit{{K}}^{\top}\right)\textit{{V}}+L\left(\textit{{V}}\right)\approx\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}}+L\left(\bar{\textit{{V}}}\right),\right.\\ &\left.g\left({\rm exp}\left(f(\tilde{\textit{{Q}}})f(\tilde{\textit{{K}}})^{\top}\right)f(\tilde{\textit{{V}}})+L\left(f(\tilde{\textit{{V}}})\right)\right)\right],\end{split}$ (16)
(a) Vanilla attention mechanism
(b) ELFATT
Figure 1: The comparison of the vanilla attention mechanism and ELFATT. For an input embedding matrix $\textit{{H}}\in\mathbb{R}^{m\times n}$, each ELFATT block will process it in two parallel attention heads to obtain two new embedding matrices, $\bar{\textit{{H}}}\in\mathbb{R}^{m\times c_{1}}$ and $\tilde{\textit{{H}}}\in\mathbb{R}^{m\times c_{2}}$, respectively. After a concatenation operation, the ELFATT block will output the updated embedding matrix $\textit{{H}}=[\bar{\textit{{H}}},\tilde{\textit{{H}}}]\in\mathbb{R}^{m\times c}$. Note: ${\rm cat}$ denotes a concatenation operation for concatenating two matrices on the column dimension into a single matrix.
Figure 2: Some visual comparison of class activation map (CAM) based attention results of different attention mechanisms obatined using Score-CAM [41]. Note: Backbone used is CSWin-T24181 [10].
Figure 3: More visual comparison of class activation map (CAM) based attention results of different attention mechanisms obatined using Score-CAM [41]. Note: Backbone used is CSWin-T24181 [10].
Eq. (12) will become as follows,

### v4

@@ Appendix A Approximation Error Bound Analysis
(13) $\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}\\ &\leq\left(\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|\textit{{U}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}+\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\right)\\ &\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}+\left\|\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
@@
If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, the total approximation error is bounded as follows,
@@
(14) $\begin{split}&\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\odot{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}\\ &\leq\left(\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}\left\|\textit{{U}}_{m}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}+\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\right)\\ &\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}+\left\|\textit{{Z}}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
For the approximation error of Eq. (9) of the main body of the paper, it is obvious to see that
$\begin{split}&\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\\ &=\left[\left({\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right)\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right)\tilde{\textit{{V}}}\right].\end{split}$
@@
(15) $\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\\ &\leq\left\|{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}-{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}\\ &+\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}-{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\geq\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, we have
@@
(16) $\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\\ &\leq\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}+\left\|\textit{{Z}}-\textit{{U}}_{m}\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
If $\left|\frac{c_{1}}{\mathscr{M}{\rm exp}\left(-0.5\right)}-1\right|<\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|$, one has
@@
(17) $\begin{split}&\left\|\left[{\rm exp}\left(\bar{\textit{{Q}}}\right){\rm exp}\left(\bar{\textit{{K}}}\right)^{\top}\bar{\textit{{V}}},\left({\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\odot\textit{{Z}}\right)\tilde{\textit{{V}}}\right]-\left[{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\bar{\textit{{V}}},{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\tilde{\textit{{V}}}\right]\right\|_{\xi}\\ &\leq\left|\frac{c_{1}}{\mathfrak{M}{\rm exp}\left(-0.5\right)}-1\right|\left\|{\rm exp}\left(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\bar{\textit{{V}}}\right\|_{\xi}+\left\|\textit{{Z}}-\textit{{U}}_{m}\right\|_{\xi}\left\|{\rm exp}\left(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top}\right)\right\|_{\xi}\left\|\tilde{\textit{{V}}}\right\|_{\xi}.\end{split}$
Inequalities (15-17) give a tighter bound than Inequalities (12-14). Fig. 2 (a) shows the upper bounds of different attention mechanisms (EFFATT (Shen et al., 2021), ELFATT, and LOCAL (Dong et al., 2022)) for approximating attention matrices of vanilla attention during the training of ImageNet-1K. The backbone used is CSWin-T (Han et al., 2023). The upper bound for ELFATT is obtained according to
@@
$\begin{split}&\left\lVert\left[\mathrm{exp}(\bar{\textit{{Q}}})\mathrm{exp}(\bar{\textit{{K}}})^{\top},\ \left(\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\odot\textit{{Z}}\right)\right]-\left[\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top}),\ \mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right]\right\rVert_{\xi}\\ &\leq\left\lVert\mathrm{exp}(\bar{\textit{{Q}}})\mathrm{exp}(\bar{\textit{{K}}})^{\top}-\mathrm{exp}(\bar{\textit{{Q}}}\bar{\textit{{K}}}^{\top})\right\rVert_{\xi}+\left\lVert\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\odot\textit{{Z}}-\mathrm{exp}(\tilde{\textit{{Q}}}\tilde{\textit{{K}}}^{\top})\right\rVert_{\xi},\end{split}$
for EFFATT is as